# L-equivalence and Fourier–Mukai partners of cubic fourfolds

### v1

@@ 3. Counting Fourier–Mukai partners of cubic fourfolds
We compute a counting formula for Fourier–Mukai partners of cubic fourfolds with certain properties. The results in this section are closely related to [HLOY02, Ma10, BBM25].
@@
Notice that the implication $(1)\Rightarrow(2)$ always holds. By [Huy17, Proposition 3.4], any equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ induces an orientation-preserving Hodge isometry $\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})\to\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ ([Huy17, Theorem 1.2]). Then, by [AT2014, Proposition 2.3], there is an identification $H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cong A_{2}^{\perp}\subset\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})$. By Voisin’s proof of the integral Hodge conjecture for cubic fourfolds, the primitive algebraic lattices coincide with the primitive Hodge classes, $A(X)_{\mathrm{prim}}=H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cap H^{2,2}(X)$. Hence we have a Hodge isometry $T(X)=A(X)_{\mathrm{prim}}^{\perp}\cong A(Y)_{\mathrm{prim}}^{\perp}=T(Y)$.
@@
If one replaces $\mathcal{H}(N)$ by $\mathcal{G}(N)$ in (3.2), one obtains the number of virtual Fourier–Mukai partners as in [BBM25].
@@
Recall from Remark 2.6 that the signature of $C_{3}$ is either $2$ or $6$ modulo $8$. In particular, we find $\operatorname{sgn}(C_{3}(-1))=-\operatorname{sgn}(C_{3})\not\equiv\operatorname{sgn}(C_{3})\pmod{8}$. Since $\operatorname{sgn}(C_{3})=\operatorname{sgn}(C_{3}^{\prime})$, we obtain $C_{3}\simeq C_{3}^{\prime}$, as required. ∎
Definition 3.8.
@@
Let $X$ be a cubic fourfold satisfying Assumption 3.1. Let $N^{\prime}$ be a lattice. Then there is a one-to-one correspondence between the set $\operatorname{FM}(X,N^{\prime})$ and the set of $O_{\text{Hodge}}(T(X))$-orbits of primitive embeddings in $\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})$.
@@
shows that the primitive embeddings $i$ and $i^{\prime}$ are in the same $O_{\text{Hodge}}(T(X))$-orbit. We will denote an embedding in this orbit by $i_{Y}$ from now on. This process thus yields a well-defined map:
$\begin{array}[]{lrcl}\operatorname{emb}:&\operatorname{FM}(X,N^{\prime})&\longrightarrow&\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})\big/O_{\text{Hodge}}(T(X))\\ &Y&\longmapsto&i_{Y}\end{array}$
The injectivity of this map follows essentially from the Torelli Theorem: if $Y,Y^{\prime}\in\operatorname{FM}(X,N^{\prime})$, such that $i_{Y}$ and $i_{Y^{\prime}}$ are in the same $O_{\text{Hodge}}(T(X))$-orbit, then there is a Hodge isometry $\psi\colon T(X)\simeq T(X)$ fitting into a commutative diagram:
@@
$\operatorname{FM}(X)=\bigsqcup_{N^{\prime}\in\mathcal{H}(N)}\operatorname{FM}(X,N^{\prime})\simeq\bigsqcup_{N^{\prime}\in\mathcal{H}(N)}\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})/O_{\text{Hodge}}(T(X)).$
@@
Let us begin by showing that $\Gamma$ is well-defined and injective. If $i,i^{\prime}\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ are two primitive embeddings contained in $\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})$. Then, by Lemma 2.4, $i$ and $i^{\prime}$ are in the same $O_{\text{Hodge}}(T(X))$-orbit, i.e. there is a commutative square
@@
where $\psi\in O_{\text{Hodge}}(T(X))$, if and only if there is an isometry $\psi^{\prime}\colon N^{\prime}\simeq N^{\prime}$ fitting into a commutative diagram
@@
This exactly means that $\gamma_{i}$ and $\gamma_{i^{\prime}}$ define the same element of $O(N^{\prime})\backslash O(A_{T(X)})/O_{\text{Hodge}}(T(X)),$ hence the map $\Gamma_{N^{\prime}}$ is well-defined and injective.
@@ 4. Examples of $\mathbb{L}$-equivalent cubic fourfolds
[IMOU20, §5] [KS18, §2.6.1] Let $X$ be a cubic fourfold with one of the following properties:
@@
In light of Proposition 4.1, Pfaffian cubic fourfolds seem like a good place to start looking for $\mathbb{L}$-equivalent cubic fourfolds. However, very general Pfaffian cubic fourfolds have no Fourier–Mukai partners:
@@
Since $X$ has an associated K3 surface $S$, we may use [PertFMPartners21, Theorem 1.1] to conclude
$\operatorname{FM}(X)=\#\operatorname{FM}(S).$
Now $S$ is a K3 surface of Picard rank 1 with degree $14$, hence [Ogu01, Proposition 1.10] implies that $\operatorname{FM}(S)=1$.
On the other hand, we may also use Theorem 3.5 directly. Since $A(X)\simeq K_{14}$ has rank $2$, it follows that $\operatorname{rk}A(X)_{\text{prim}}=1$. Therefore, $N\coloneqq A(X)_{\text{prim}}$ is clearly unique in its genus. Moreover, we have $\operatorname{disc}(T(X))=14$, and $A_{T(X)}$ is cyclic, since it is naturally a subgroup of the cyclic group $A_{N}$ by Remark 3.4. This implies that $A_{T(X)}$ has precisely $2$ isometries, namely $\pm\operatorname{id}_{A_{T(X)}}$. Indeed, if $x\in A_{T(X)}$ is a generator, then for any $\sigma\in O(A_{T(X)})$, we must have $\sigma(x)=ax$ for some $a\in(\mathbb{Z}/14\mathbb{Z})^{\times}$. However, since $x^{2}=\sigma(x)^{2}=a^{2}x^{2}\pmod{2\mathbb{Z}}$, it follows that we must have $a^{2}\equiv 1\pmod{14}$. Since $\pm 1$ are the only elements of $(\mathbb{Z}/14\mathbb{Z})^{\times}$ whose square equals $1$, we find $O(A_{T(X)})=\{\pm\operatorname{id}_{A_{T(X)}}\}$. Since both of these isometries are in the image of the group homomorphism $O_{\text{Hodge}}(T(X))\to O(A_{T(X)})$, it follows from Theorem 3.5 that $\operatorname{FM}(X)=1$. ∎
@@
Some of the lattice theoretical computations involving the next examples were done with OSCAR [OSCAR, OSCAR-book]. We briefly recall here the behaviour of the key functions we used.
  • primitive_embeddings(L, N) (developed by Stevell Muller): computes primitive embeddings of the lattice L into N (optionally with classification data such as orthogonal complements).
@@
Since $\operatorname{disc}(T(X))=\operatorname{disc}(A(X))=21$ is divisible by $3$, Theorem 3.5 is not applicable in this situation. Therefore, we use the counting algorithm of [BBM25].
@@
$\operatorname{FM}(X)=|O(N)\backslash O(A_{T(X)})/O_{\text{Hodge}}(T(X))|,$
@@
where $n\geq 1$. If $X$ is a cubic fourfold with $A(X)\simeq L_{n}$, then $X$ satisfies property B of Proposition 4.1, so we are able to compute the class of $X$ in $K_{0}(\operatorname{Var}_{\mathbb{C}})$. By Proposition 4.4, $X$ has no Fourier–Mukai partners when $n=3$. Also, since the discriminant of $L_{n}$ is $\operatorname{disc}(L_{n})=8n-3$, we may use Theorem 3.5 to compute the number of Fourier–Mukai partners of $X$ whenever $n$ is not divisible by $3$.
@@
Let $X$ be a cubic fourfold such that $A(X)\simeq L_{10}$. Assume moreover that $O_{\text{Hodge}}(T(X))\simeq\mathbb{Z}/2\mathbb{Z}$. Then $\operatorname{FM}(X)=2$. Moreover, the unique non-trivial Fourier–Mukai partner $Y$ of $X$ has $A(Y)_{\text{prim}}\simeq A(X)_{\text{prim}}$. In addition, $X$ and $Y$ are trivially $\mathbb{L}$-equivalent.
@@
$|O(N)\backslash O(A_{T(X)})/O_{\text{Hodge}}(T(X))|=2.$
@@
and one can compute using [HLOY02, Theorem 2.3] that $S$ has no non-trivial Fourier–Mukai partner, hence $S\simeq S^{\prime}$. In conclusion, we have

### v2

@@ 3. Counting Fourier–Mukai partners of cubic fourfolds
We compute a counting formula for Fourier–Mukai partners of cubic fourfolds with certain properties. The results in this section are closely related to [19, 34, 50, 13, 9].
@@
Notice that the implication $(1)\Rightarrow(2)$ always holds. By [23, Proposition 3.4], any equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ induces an orientation-preserving Hodge isometry $\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})\to\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ ([23, Theorem 1.2]). Then, by [2, Proposition 2.3], there is an identification $H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cong A_{2}^{\perp}\subset\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})$. By Voisin’s proof of the integral Hodge conjecture for cubic fourfolds, the primitive algebraic lattices coincide with the primitive Hodge classes, $A(X)_{\mathrm{prim}}=H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cap H^{2,2}(X)$. Hence we have a Hodge isometry $T(X)=A(X)_{\mathrm{prim}}^{\perp}\cong A(Y)_{\mathrm{prim}}^{\perp}=T(Y)$.
@@
On the other hand, if $\operatorname{disc}(N)$ is not divisible by $3$, by the symmetry of the situation we obtain an action of $O(T)$ on $O(A_{N})$.
@@
If one replaces $\mathcal{H}(N)$ by $\mathcal{G}(N)$ in (3.2), one obtains the number of virtual Fourier–Mukai partners as in [9].
Definition 3.7.
@@
Recall from Remark 2.6 that the signature of $C_{3}$ is either $2$ or $6$ modulo $8$. In particular, we find $\operatorname{sgn}(C_{3}(-1))=-\operatorname{sgn}(C_{3})\not\equiv\operatorname{sgn}(C_{3})\pmod{8}$. Since $\operatorname{sgn}(C_{3})=\operatorname{sgn}(C_{3}^{\prime})$, we obtain $C_{3}\simeq C_{3}^{\prime}$, which shows that $N$ and $N^{\prime}$ are in the same genus.
We now check that $N^{\prime}\in\mathcal{H}(N)$. Firstly, note that if $N^{\prime}$ were to contain a vector of square $2$, this would be a contradiction with the surjectivity of the period map, see Theorem 2.10.
Similarly, we now prove that $N^{\prime}$ cannot contain a vector of square $6$ and divisibility $3$. For the sake of contradiction, let us suppose that there exists a vector $v\in N^{\prime}$ with $v^{2}=6$ and $\operatorname{div}_{N^{\prime}}(v)=3$. Then the element $\frac{v}{3}\in A_{N^{\prime}}\simeq A_{T(Y)(-1)}\oplus C_{3}$ has order $3$, hence it is a generator for $C_{3}$ and, importantly, is not contained in $A_{T(Y)(-1)}$. The gluing group $G\coloneqq\frac{H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}}{N\oplus T(Y)}$ is equal to $A_{T(Y)(-1)}\oplus A_{T(Y)}$ as a subgroup of $A_{N^{\prime}}\oplus A_{T(Y)}$. Therefore, we have a natural isomorphism $A_{H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}}\simeq G^{\perp}/G\simeq C_{3}$, showing that $A_{H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}}$ is also generated by $\frac{v}{3}$. This implies that $\frac{v}{3}\in H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}^{\vee}$, hence $v$ must have divisibility $3$ in $H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}$. This contradicts the surjectivity of the period map, c.f. Theorem 2.10.
This shows that $N^{\prime}\in\mathcal{H}(N)$, as required. ∎
@@
Let $X$ be a cubic fourfold satisfying Assumption 3.1. Let $N^{\prime}$ be a lattice. Then there is a one-to-one correspondence between the set $\operatorname{FM}(X,N^{\prime})$ and the set of $O_{\mathrm{Hodge}}(T(X))$-orbits of primitive embeddings in $\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})$.
@@
shows that the primitive embeddings $i$ and $i^{\prime}$ are in the same $O_{\mathrm{Hodge}}(T(X))$-orbit. We will denote an embedding in this orbit by $i_{Y}$ from now on. This process thus yields a well-defined map:
$\begin{array}[]{lrcl}\operatorname{emb}:&\operatorname{FM}(X,N^{\prime})&\longrightarrow&\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})\big/O_{\mathrm{Hodge}}(T(X))\\ &Y&\longmapsto&i_{Y}\end{array}$
The injectivity of this map follows essentially from the Torelli Theorem: if $Y,Y^{\prime}\in\operatorname{FM}(X,N^{\prime})$, such that $i_{Y}$ and $i_{Y^{\prime}}$ are in the same $O_{\mathrm{Hodge}}(T(X))$-orbit, then there is a Hodge isometry $\psi\colon T(X)\simeq T(X)$ fitting into a commutative diagram:
@@
$\operatorname{FM}(X)=\bigsqcup_{N^{\prime}\in\mathcal{H}(N)}\operatorname{FM}(X,N^{\prime})\simeq\bigsqcup_{N^{\prime}\in\mathcal{H}(N)}\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})/O_{\mathrm{Hodge}}(T(X)).$
@@
Let us begin by showing that $\Gamma$ is well-defined and injective. If $i,i^{\prime}\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ are two primitive embeddings contained in $\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})$. Then, by Lemma 2.4, $i$ and $i^{\prime}$ are in the same $O_{\mathrm{Hodge}}(T(X))$-orbit, i.e. there is a commutative square
@@
where $\psi\in O_{\mathrm{Hodge}}(T(X))$, if and only if there is an isometry $\psi^{\prime}\colon N^{\prime}\simeq N^{\prime}$ fitting into a commutative diagram
@@
This exactly means that $\gamma_{i}$ and $\gamma_{i^{\prime}}$ define the same element of $O(N^{\prime})\backslash O(A_{T(X)})/O_{\mathrm{Hodge}}(T(X)),$ hence the map $\Gamma_{N^{\prime}}$ is well-defined and injective.
@@
Due to the symmetry of the situation, one might expect an similar counting formula for Fourier–Mukai partners of cubic fourfolds to hold when we assume $\operatorname{disc}(A(X)_{\mathrm{prim}})$, rather than $\operatorname{disc}(T(X))$, is not divisible by $3$. The fundamental difference between these two assumptions is the following. If $X$ is a cubic fourfold such that $\operatorname{disc}(T(X))$ is not divisible by $3$, then $\operatorname{disc}(T(Y))$ is also not divisible by $3$ for any Fourier–Mukai partner $Y$ of $X$. On the other hand, if we assume $\operatorname{disc}(A(X)_{\mathrm{prim}})$ to be indivisible by $3$, then $X$ may have several Fourier–Mukai partners $Y$ for which $\operatorname{disc}(A(Y)_{\mathrm{prim}})$ is a multiple of $3$. Nevertheless, we have the following Counting Formula for Fourier–Mukai partners of $X$ whose primitive algebraic lattice has discriminant not divisible by $3$.
Proposition 3.12.
Let $X$ be a cubic fourfold satisfying Assumption 3.1, and write $N\coloneqq A(X)_{\mathrm{prim}}$. Suppose that $\operatorname{disc}(N)$ is not divisible by $3$. Then for any $N^{\prime}\in\mathcal{H}(N)$, we have
(3.3) $\operatorname{FM}(X,N^{\prime})=|O(N^{\prime})\backslash O(A_{N})/O_{\mathrm{Hodge}}(T(X))|.$
Here, the action of $O_{\mathrm{Hodge}}(T(X))$ on $O(A_{N})$ is the one defined by Remark 3.4.
Proof.
We fix an isometry $A_{T(X)}\simeq A_{N^{\prime}}\oplus C_{3}$, where $C_{3}$ is one of the two non-degenerate finite quadratic modules of order $3$. A primitive embedding $i\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ induces a pair $(G_{i},\gamma_{i})$, where $G_{i}\subset A_{T(X)}$ and $\gamma_{i}\colon G_{i}\hookrightarrow A_{N^{\prime}}$. The group $G_{i}$ is equal to $A_{N^{\prime}}$ as a subgroup of $A_{T(X)}$, since $A_{N^{\prime}}\subset A_{T(X)}$ is the unique maximal subgroup whose order is not divisible by $3$. Therefore, $\gamma_{i}$ can be viewed as an element of $O(A_{N^{\prime}}).$
Similarly to the proof of Theorem 3.5, we obtain a map
[…]
which we claim to be a bijection. The proof the bijectivity of $\Gamma_{N^{\prime}}$ is completely analogous to the proof of Theorem 3.5, so we omit it here. ∎
Remark 3.13.
In the setting of Proposition 3.12, it is possible that $X$ has a Fourier–Mukai partner $Y$ with $N^{\prime}=A(Y)_{\mathrm{prim}}\notin\mathcal{H}(N)$. This is equivalent to the discriminant of $A(Y)_{\mathrm{prim}}$ being divisible by $3$. In this case, the gluing group of the embedding $N^{\prime}\hookrightarrow H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}$ is isomorphic to $A_{T(Y)}$, but it is no longer necessarily the case that $A_{N^{\prime}}$ can be written as $A_{T(Y)}\oplus C_{3}$ for some finite quadratic form $C_{3}$ of order $3$. It is still possible to compute $\operatorname{FM}(X,N^{\prime})$ via Proposition 3.10, c.f. [9], but a simple formula such as (3.3) seems to be unattainable in this case. See Remark 4.8 for examples of this phenomenon.
@@ 4. Examples of $\mathbb{L}$-equivalent cubic fourfolds
[25, §5] [30, §2.6.1] Let $X$ be a cubic fourfold with one of the following properties:
@@
In light of Proposition 4.1, Pfaffian cubic fourfolds seem like a good place to start looking for $\mathbb{L}$-equivalent cubic fourfolds. However, very general Pfaffian cubic fourfolds have no Fourier–Mukai partners. This result follows immediately from [13, Proposition 2.6], and also from [50, Theorem 1.1] combined with [45, Proposition 1.10]. Here, we give a proof using Theorem 3.5.
@@
Since $A(X)\simeq K_{14}$ has rank $2$, it follows that $\operatorname{rk}A(X)_{\text{prim}}=1$. Therefore, $N\coloneqq A(X)_{\text{prim}}$ is clearly unique in its genus. Moreover, we have $\operatorname{disc}(T(X))=14$, and $A_{T(X)}$ is cyclic, since it is naturally a subgroup of the cyclic group $A_{N}$ by Remark 3.4. This implies that $A_{T(X)}$ has precisely $2$ isometries, namely $\pm\operatorname{id}_{A_{T(X)}}$. Indeed, if $x\in A_{T(X)}$ is a generator, then for any $\sigma\in O(A_{T(X)})$, we must have $\sigma(x)=ax$ for some $a\in(\mathbb{Z}/14\mathbb{Z})^{\times}$. However, since $x^{2}=\sigma(x)^{2}=a^{2}x^{2}\pmod{2\mathbb{Z}}$, it follows that we must have $a^{2}\equiv 1\pmod{14}$. Since $\pm 1$ are the only elements of $(\mathbb{Z}/14\mathbb{Z})^{\times}$ whose square equals $1$, we find $O(A_{T(X)})=\{\pm\operatorname{id}_{A_{T(X)}}\}$. Since both of these isometries are in the image of the group homomorphism $O_{\mathrm{Hodge}}(T(X))\to O(A_{T(X)})$, it follows from Theorem 3.5 that $\operatorname{FM}(X)=1$. ∎
@@
Some of the lattice theoretical computations involving the next examples were done with OSCAR [48, 10]. We briefly recall here the behaviour of the key functions we used.
  • primitive_embeddings(L, N) (developed by Stevell Muller): computes primitive embeddings of the lattice N into L (optionally with classification data such as orthogonal complements).
@@
Since $\operatorname{disc}(T(X))=\operatorname{disc}(A(X))=21$ is divisible by $3$, Theorem 3.5 is not applicable in this situation. Therefore, we use the counting algorithm of [9].
@@
$\operatorname{FM}(X)=|O(N)\backslash O(A_{T(X)})/O_{\mathrm{Hodge}}(T(X))|,$
@@
where $n\geq 1$. If $X$ is a cubic fourfold with $A(X)\simeq L_{n}$, then $X$ satisfies property B of Proposition 4.1, so we are able to compute the class of $X$ in $K_{0}(\operatorname{Var}_{\mathbb{C}})$.
Importantly, if $(S_{1}$, $S_{2}$, $S_{3})$ is an ordered $\mathbb{Z}$-basis for $A(X)$ whose associated Gram matrix is $L_{n}$, then $S_{1}$ is the square of a hyperplane class if and only if $n$ is odd. On the other hand, if $n$ is even, then $S_{2}$ is the square of a hyperplane class.
By Proposition 4.4, $X$ has no Fourier–Mukai partners when $n=3$. Also, since the discriminant of $L_{n}$ is $\operatorname{disc}(L_{n})=8n-3$, we may use Theorem 3.5 to compute the number of Fourier–Mukai partners of $X$ whenever $n$ is not divisible by $3$.
@@
Let $X$ be a cubic fourfold such that $A(X)\simeq L_{10}$. Assume moreover that $O_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}/2\mathbb{Z}$. Then $\operatorname{FM}(X)=2$. Moreover, the unique non-trivial Fourier–Mukai partner $Y$ of $X$ has $A(Y)_{\text{prim}}\simeq A(X)_{\text{prim}}$. In addition, $X$ and $Y$ are trivially $\mathbb{L}$-equivalent.
@@
$|O(N)\backslash O(A_{T(X)})/O_{\mathrm{Hodge}}(T(X))|=2.$
@@
and one can compute using [19, Theorem 2.3] that $S$ has no non-trivial Fourier–Mukai partner, hence $S\simeq S^{\prime}$. In conclusion, we have
@@
We now turn our attention to Pfaffian cubic fourfolds, and consider the lattice
$L_{a,b}\coloneqq\left(\begin{matrix}14&a\\ a&2b\end{matrix}\right),$
where $a,b\in\mathbb{Z}$ are chosen in such a way that $L_{a,b}$ is a non-degenerate positive-definite lattice. We will also write $A_{a,b}\coloneqq A_{L_{a,b}}.$
Proposition 4.7.
Let $X$ be a cubic fourfold such that $A(X)\simeq L_{2,9}$. Assume moreover that $O_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}/2\mathbb{Z}$. Then $\operatorname{FM}(X,L_{2,9})=2$. That is, $X$ has a unique Fourier–Mukai partner $Y$ with the property that $A(Y)_{\mathrm{prim}}\simeq A(X)_{\mathrm{prim}}$. Moreover, $X$ and $Y$ are trivially $\mathbb{L}$-equivalent.
Proof.
Since $\operatorname{disc}(L_{2,9})=248$ is not divisible by $3$, we may use Proposition 3.12 to compute
$\operatorname{FM}(X,L_{2,9})=|O(L_{2,9})\backslash O(A_{2,9})/O_{\mathrm{Hodge}}(T(X))|.$
We find $O(A_{2,9})\simeq\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2\mathbb{Z}$ and $O(L_{2,9})\simeq\mathbb{Z}/2\mathbb{Z}$. Thus, since $O_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}/2\mathbb{Z}$ by assumption, we obtain
$\operatorname{FM}(X,L_{2,9})=2.$
Denote by $Y$ the unique Fourier–Mukai partner of $X$ with $A(Y)_{\mathrm{prim}}\simeq L_{2,9}$. Then we have $Y\in\mathcal{C}_{14}$, i.e. $Y$ is a Pfaffian cubic fourfold and thus satisfies Property A of Proposition 4.1. In particular, we find K3 surfaces $S$, $S^{\prime}$ such that
Moreover, $S$ and $S^{\prime}$ are Fourier–Mukai partners. One can compute that
$\operatorname{NS}(S)\simeq\left(\begin{matrix}26&22\\ 22&-10\end{matrix}\right),$
but such a K3 surface never has non-trivial Fourier–Mukai partners, so we have $S\simeq S^{\prime}$, hence $X$ and $Y$ are trivially $\mathbb{L}$-equivalent. ∎
Remark 4.8.
A cubic fourfold $X$ as in Proposition 4.7 can be shown to have more Fourier–Mukai partners. For example, the genus of $L_{2,9}$ consists of three lattices:
$L_{2,9}=\left(\begin{matrix}14&2\\ 2&18\end{matrix}\right),\quad N_{1}\coloneqq\left(\begin{matrix}126&16\\ 16&4\end{matrix}\right)\simeq\left(\begin{matrix}62&0\\ 0&4\end{matrix}\right),\quad N_{2}\coloneqq\left(\begin{matrix}126&2\\ 2&2\end{matrix}\right).$
However, $N_{2}$ clearly contains a vector of square $2$, hence $N_{2}\notin\mathcal{H}(L_{2,9})$, and $X$ has no Fourier–Mukai partner $X^{\prime}$ with $A(X^{\prime})_{\mathrm{prim}}\simeq N_{2}$. On the other hand, $N_{1}$ contains neither vectors of square 2 nor vectors of square 6 and divisibility 3. One computes using Proposition 3.12 that $X$ has precisely $1$ Fourier–Mukai partner $X^{\prime}$ with $A(X^{\prime})_{\mathrm{prim}}\simeq N_{1}$.
Moreover, there is a primitive embedding $T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ whose orthogonal complement is isometric to the lattice
$N_{3}\coloneqq\left(\begin{matrix}66&12\\ 12&36\end{matrix}\right).$
Importantly, $N_{3}$ has discriminant $2242$, which is divisible by $3$. Since $N_{3}$ contains no vectors of square $6$ nor vectors of square $2$, it follows that $X$ has at least one Fourier–Mukai partner $X^{\prime}$ with $A(X^{\prime})_{\mathrm{prim}}\simeq N_{3}$. Finally, the genus of $N_{3}$ consists of $N_{3}$ and the lattice
$N_{4}\coloneqq\left(\begin{matrix}18&-60\\ -60&324\end{matrix}\right),$
which again contains no vectors of square $6$ or $2$, hence $X$ has at least one Fourier–Mukai partner corresponding to $N_{4}$ as well.
It is an interesting question whether the Fourier–Mukai partners of $X$ whose primitive algebraic lattices are not isometric to that of $X$ are $\mathbb{L}$-equivalent to $X$.
Remark 4.9.
The techniques of the proof of Proposition 4.7 can be used to prove that the next lattices which generate examples of trivially $\mathbb{L}$-equivalent Pfaffian cubic fourfolds in precisely the same way as Proposition 4.7 are
$L_{4,6},\quad L_{5,6},\quad L_{6,5},\quad L_{6,10},\quad L_{8,6},\quad L_{9,8}.$